Find two consecutive intergers such that if 9 times the first is subtracted from 5 times the second, then the result is 133.
Answer by macston(5194)

.
F=first; S=second=F+1
.
5S-9F=133 Substitute for S.
5(F+1)-9F=133
5F+5-9F=133
-4F=128
F=-32 ANSWER 1: The first is -32.
S=F+1=-32+1=-31 ANSWER 2: the second is -31.
.
CHECK:
5S-9F=133
5(-31)-9(-32)=133
-155+288=133
133=133